40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1040     Entry time: Fri Oct 10 13:57:33 2008
Author: Alberto 
Type: Omnistructure 
Category: Computers 
Subject: Problems in locking the X arm 
This morning for some reason that I didn't clearly understand I could not lock the Xarm. The Y arm was not a problem and the Restore and Align script worked fine.

Looking at the LSC medm screen something strange was happening on the ETMX output. Even if the Input switch for c1:LSC-ETMX_INMON was open, there still was some random output going into c1:LSC-ETMX_INMON, and it was not a residual of the restor script running. Probably something bad happened this monring when we rebooted all the FE computers for the RFM network crash that we had last night.

Restarting the LSC computer didn't solve the problem so I decided to reboot the scipe25 computer, corresponding to c1dcuepics, that controls the LSC channels.

Somehow rebooting that machine erased all the parameters on almost all medm screens. In particular the mode cleaner mirrors got a kick and took a while to stop. I then burtrestored all the medm screen parameters to yesterday Thursday October 9 at 16:00. After that everything came back to normal. I had to re-lock the PMC and the MC.

Burtrestoring c1dcuepics.snap required to edit the .snap file because of a bug in burtrestore for that computer wich adds an extra return before the final quote symbol in the file. That bug should be fixed sometime.

The rebooting apparently fixed the problem with ETMX on the LSC screen. The strange output is not present anymore and I was able to easily lock the X arm. I then run the Align and the Restore full IFO scripts.
ELOG V3.1.3-